Systems Staff Meeting Notes 4/2/01

Systems Staff Meeting Notes
April 2, 2001

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Systems staff announcement: Bobbie Meng is leaving for a job in Tigert Hall. Her last day is April 13, 2001.
  2. Andy Zahn is no longer working for us. Debra will take over as mistress of procurement.
Review Minutes of 3/26/01
  1. Ann #2 -Annoying beeps from CLAS server probes – CLAS has taken care of the problem.
  2. Nets #6 – Personnel has not picked up their laptop yet.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. New Web server incoming – Compaq: it has four processors, 2MB, 2 p/s, 180 GIG as well as a very big battery.
  2. 32 laptops for the new classroom in MSL - $3000 Dells (1 GIG processors with W2K in order to be workstation equivalent) – order is on hold, it’s not as sure of a thing as the Web server.
  3. Ongoing issue with Dell regarding the video problem with a number of machines – Dell shipped 10 dongles on Friday. They will probably actually send ferrite beads, not dongles. They are supposed to help with flickering monitor problem.
  4. A Dell technician came & fixed a dead display on one of the laptops that is destined for one of the departments.
Liaisons
  1. New patch for IE – Bill put out an update late Friday afternoon. It is a critical update – targeting HTML. If the download goes out but stops at some point & doesn’t finish – it is probably hitting Microsoft during a busy period. If it starts to download and stops partway through, it's on the Microsoft end.
Networking
  1. Weekend Power Outage – Bill & Will came in on Saturday to assist – power went down at around 7:20 AM. Turns out, they really only needed the power in part of the North Tower to go down. Our server circuits stayed hot – although we went through the trouble of powering them down. Due to this kind of thing, we’ll be pushing to get the building electrical mapped out. The Web & squid servers remained up the whole time. The Xerox print server lost power. The Novell print server would not come up. Finally, it magically booted up after many power-ons. The Web forms server was still off this morning – just needed to turn it back on.
  2. Building Security System – apparently didn’t convert to daylight savings time. Some employees could not gain access. Seems to be resolved now.
  3. Telephones in the room 229 West closet – one is to be used and one not – it is said that it can take down the system if used. Need one of the phones to have outside access and a phone number that can be called to. Really, all of the closets should have functioning phone access. Need to check & verify.
  4. SID will be replaced with Plug-in.
  5. Room 100 – desk has been ordered for the Circulation point.
  6. MSL 215 – scheduled for wiring on Thursday, if the furniture is assembled.
  7. Hole in the wall should be plugged in. Wiring walk-thru for 4th-6th floors in West was done with Tim.
  8. MSL 107 conference room – 32 drops? Not decided yet, money uncertain.
  9. Ports in stacks – West wires in ceiling. Gary Cornwell has requested wireless antennas. He said he would submit the paperwork, but no one has seen it yet. Will need at least 2 access points per floor. Could be pricey due to signal blockage. May put them on the 3rd & 5th floors. Will have to conduct tests.
  10. West – closet on the 5th floor – rewiring to be able to get rid of the hub on the 4th floor that feeds the 5th & 6th floor. Do we want to put a fiber switch on the 5th floor? Yes.
  11. Per Christian Conrad (UF Webmaster), a Web camera will be placed outside the 5th floor of West, filming out onto the Plaza of Americas.
Software Projects Status
  1. Preservation Binding – Hilton has finished.
  2. LASSIE modifications – Hilton has finished.
  3. OPS Tracker for Resource Services – Hilton is currently working on it.
  4. Personnel database mods – ongoing.
  5. Cynthia is still working on the Elected Officials database.
  6. Bill & Winston will be conducting a comprehensive review of all Systems-supported software projects this month.
  7. DLC – 1st version of Ringling is complete and ready for review on Wednesday. It also uses some SID images – the user has to download the plug-in.
DP Coordinator
  1. Business tapes are almost completely removed and destroyed from the Tape Library. Next step will be the actual physical movement of the remaining tapes. Rack/slots will have to be reassigned. LTMS software will need to be modified. The power for the server room should be complete by April 28 (which just happens to be Will’s birthday).


Back to the Systems Homepage

Last updated April 9, 2001
by Debra Harris
debharr@mail.uflib.ufl.edu